OpenStudy (anonymous):

What is the area of a circle with a diameter of 22?

hartnn (hartnn):

\(Area ~ of ~circle = \pi r^2 = \pi d^2/4\) d = diameter = 22

OpenStudy (anonymous):

379.94^2 Is that right?

hartnn (hartnn):

yes, i got around 380 too

hartnn (hartnn):

wait, whats that ^2 ?

OpenStudy (anonymous):

Squared, I thought you square when talking about area.

hartnn (hartnn):

A = pi* 22^2/4 = 380.13 units^2

OpenStudy (anonymous):

OK, thanks for telling me.

hartnn (hartnn):

the units are squared, like for example if the diamamter was 22 cm, then area = 380.13 cm^2

hartnn (hartnn):

welcome ^_^